Camaro GTO Torino
Cash $ 1,900 $ 140 $ 1,000
Short-term investments 0 0 700
Current receivables 170 420 500
Inventory 1,950 1,840 2,600
Prepaid expenses 200 500 900
Total current assets $ 4,220 $ 2,900 $ 5,700
Current liabilities $ 2,040 $ 1,230 $ 3,100


Compute the current ratio and acid-test ratio for each of the following separate cases.
Answer #1

Current ratio = Total current assets / Current liabilities

Camaro = 4220/2040 = 2.07

GTO = 2900/1230 = 2.36

Torino = 5700/3100 = 1.84

Acid-test ratio = (Total current assets - Inventory - Prepaid expenses)

Camaro = (4220-1950-200)/2040 = 1.01

GTO = (2900-1840-500)/1230 = 0.46

Torino = (5700-2600-900)/3100 = 0.71